# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

### v1

@@ 5.2 Comparison with Gram–Schmidt Random Walk Design
harshaw2022design propose an alternative, design-based approach to variance reduction. Rather than modifying the estimator, they optimize the randomization mechanism itself through the Gram–Schmidt walk (GSW) design, which ensures covariate balance while maintaining marginal assignment probabilities. While such algorithmic designs can achieve improved covariate balance in finite samples, they require using the covariates prior to the experiment when constructing the design. In contrast, under simple or complete random assignment, one can freely add or remove covariates, analyze different subgroups of the population, or draw additional samples even after the experiment has been conducted. This flexibility—together with the transparency and ease of implementation of standard randomization— makes simple and complete random designs more attractive in most empirical settings.
harshaw2022design show that their GSW design achieves the following variance bound for any choice of $\phi\in(0,1)$:
$\frac{1}{n^{2}}\min_{\boldsymbol{\beta}\in\mathbb{R}^{d}}\Big[\frac{1}{\phi}\left\|{\boldsymbol{\mu}}-{\bm{X}}\boldsymbol{\beta}\right\|_{2}^{2}+\frac{\zeta_{{\bm{X}}}^{2}}{1-\phi}\left\|\boldsymbol{\beta}\right\|_{2}^{2}\Big].$ (18)
We now compare this bound with that of our LOORA-HT estimator.
The second term in (18) and the corresponding regularization term in (7) are not directly comparable in general, although both scale on the order of $k/n^{2}$. Setting $\lambda=\zeta_{{\bm{X}}}^{2}/(1-\phi)$, Lemma˜2 implies that
$h_{i}(\widetilde{{\bm{X}}},\tfrac{\zeta_{\widetilde{{\bm{X}}}}^{2}}{1-\phi})\;\leq\;\frac{1-\phi}{2-\phi}.$
@@
$\frac{1}{1-\big(h_{i}(\widetilde{{\bm{X}}},\tfrac{\zeta_{\widetilde{{\bm{X}}}}^{2}}{1-\phi})\big)^{2}}\;\leq\;(2-\phi)^{2}.$
For $\phi\in[0,(3-\sqrt{5})/2)$, we have $(2-\phi)^{2}\leq 1/\phi$. Hence, the scaling factor that appears in our variance bound is uniformly smaller than that of the GSW design. Moreover, the first term in (18) diverges as $\phi\to 0$, whereas the corresponding constant in the variance of LOORA-HT remains bounded by $4$.
Finally, when the leverage scores of $\widetilde{{\bm{X}}}$ are small—that is, $h_{i}(\widetilde{{\bm{X}}},0)$ is small for all ii—we may safely take $\lambda=0$ in our estimator. In contrast, for the GSW design, the smallest feasible regularization parameter is $\lambda=\zeta_{{\bm{X}}}^{2}$, so the variance bound cannot achieve the same limiting rate. This highlights a key practical advantage of LOORA-HT under standard randomization schemes: it achieves comparable or tighter variance guarantees without altering the experimental design itself.
5.3 Connections of LOORA-DM to Leave-Two-Out Estimators
spiess2025optimal provides a general characterization of unbiased design-based estimators for average treatment effects under arbitrary randomization schemes. He shows that for Horvitz–Thompson (HT) and difference-in-means (DM) type estimators, unbiasedness imposes specific structural constraints on how each unit’s outcome can depend on treatment assignments of other units. In particular, the HT estimator must have a leave-one-out structure—each adjusted outcome for unit $i$ may depend on the treatment assignments of all other units except ii—whereas the DM estimator must satisfy a stronger leave-two-out condition, meaning that each pairwise contribution between units $i$ and $j$ can depend only on the treatment assignments of units other than $i$ and $j$.
In what follows, we show that the LOORA-DM estimator satisfies exactly this leave-two-out property, and thus conforms to the necessary unbiasedness conditions established by spiess2025optimal.
We show that LOORA-DM estimator can be written as
@@ 5.3 Connections of LOORA-DM to Leave-Two-Out Estimators
where ${\bm{z}}_{-ij}$ denotes the treatment assignments of all units except $i$ and $j$. Note that
@@
Adjustment for treated and control units.
where
For a control unit $e$ ($d_{e}=0$), the corresponding adjustment is
with
@@
Define a unified adjustment quantity
$\widetilde{y}_{\ell}=\begin{cases}\frac{n_{C}(n-1)}{(n_{T}-1)n}y^{(1)}_{\ell},&d_{\ell}=1,\\[2.58334pt] \frac{n_{T}(n-1)}{(n_{C}-1)n}y^{(0)}_{\ell},&d_{\ell}=0,\end{cases}$
@@
Simple algebra then shows that, for treated units ($d_{i}=1$), the right-hand side reduces exactly to (19), and for control units ($d_{i}=0$) to (20). Thus, $\widehat{\tau}_{\textup{LDM}}$ can indeed be expressed in the leave-two-out form above.
Interpretation.
This representation shows that LOORA-DM satisfies the structural unbiasedness conditions identified by spiess2025optimal. Specifically, the adjustment $\phi_{ij}({\bm{z}}_{-ij})$ for each pair $(i,j)$ depends only on the treatment assignments of the remaining units, ensuring that the contribution of each pair is conditionally unbiased given the rest of the assignment vector. Hence, while the LOORA-HT estimator satisfies the leave-one-out property required for Horvitz–Thompson estimators, LOORA-DM naturally satisfies the stronger leave-two-out property required for difference-in-means estimators. Our construction therefore provides a concrete, closed-form realization of the abstract unbiasedness conditions derived in spiess2025optimal.
6 Experiments
This section provides empirical evidence on the finite-sample performance of our proposed estimators. We use experimental data from two within-subject studies, which allow us to observe both treatment and control outcomes for each unit. These data make it possible to generate synthetic randomized experiments under controlled conditions, enabling a systematic comparison of estimator bias, efficiency, and confidence-interval coverage. Our goal is not to assess the substantive findings of the original studies, but rather to evaluate the statistical behavior of LOORA–HT and LOORA–DM relative to standard regression-adjusted estimators.
The datasets are drawn from allcott2015evaluating and mcdonald2025evaluating. A brief description of each is provided below, and our replication code—available online—contains full details for data processing and experimental design.
@@ 6 Experiments
    Statehood dataset. mcdonald2025evaluating study the persuasiveness of policy arguments using a repeated-measures, within-subject design. Respondents first record their opinions on a policy issue, then read one or more arguments either supporting or opposing the policy, and finally re-state their opinions after exposure to each argument. We use the portion of their data concerning opinions about granting statehood to the District of Columbia (DC). The outcome under control is the respondent’s opinion before reading any arguments; the outcome under treatment is the opinion after reading (i) an argument against DC statehood emphasizing political corruption, and (ii) an argument in favor highlighting taxation without representation. We restrict attention to male respondents aged 40–49. Covariates include one-hot encodings of six categorical variables: party affiliation, race, voter registration, political attentiveness, education, and ideology. The resulting dataset contains $36$ units and $32$ covariates.
@@
    Lightbulb dataset. allcott2015evaluating conduct an experiment on consumer preferences between two types of lightbulbs offered at varying relative prices. Respondents first choose between the two options, after which they receive information about the energy costs of each type and make the choice again. The initial choice forms the control outcome; the post-information choice forms the treatment outcome. Covariates are obtained by one-hot encoding categorical features from the original data, including employment status, renter status, U.S. region, metropolitan statistical area status, marital status, income, housing type, gender, and ethnicity. The final dataset contains $123$ units and $53$ covariates.
Dataset Method Bias STD RMSE CI coverage (95%) CI average length (95%)
Statehood REG 0.0504 0.5191 0.5216 0.7182 0.5783
Ridge-REG 0.0887 0.2708 0.2849 0.8956 0.4467
LOORA-HT 0.0010 0.3125 0.3125 0.9425 0.6274
LOORA-DM 0.0510 0.2823 0.2869 0.9461 0.5880
Lightbulb REG 0.0021 0.1054 0.1054 0.9115 0.1797
Ridge-Reg 0.0274 0.0855 0.0898 0.9198 0.1542
LOORA-HT 0.0000 0.0849 0.0849 0.9583 0.1746
LOORA-DM 0.0021 0.0799 0.0799 0.9619 0.1668
Table 2: Performance metrics for four methods on two datasets with treatment assignment probabilities correlated with covariates.
Because both datasets arise from within-subject designs, we observe for each individual both potential outcomes. This feature allows us to simulate many synthetic randomized experiments by masking one outcome per unit according to a prescribed randomization rule. We can then evaluate the finite-sample properties of each estimator—bias, standard deviation, root mean squared error (RMSE), and the average length and coverage of confidence intervals—under realistic covariate distributions.
We consider two treatment-assignment mechanisms. In both cases, assignments are independent across units. In the first design, the treatment probability varies with the covariates. Specifically, we draw a random Gaussian vector, compute for each unit $i$ the cosine similarity $c_{i}$ between this vector and its covariate vector, and set
@@
In the second design, the treatment probability is fixed at $p_{i}=0.5$ for all units.
Each synthetic experiment is repeated 100,000100{,}000 times. Tables 2 and 3 summarize the results. We compare our estimators, LOORA–HT and LOORA–DM, with two benchmarks: (i) the classical regression-adjusted estimator (REG) of lin2013agnostic, and (ii) a ridge-regularized variant (Ridge–REG).
Our estimators consistently achieve lower RMSE than REG, reflecting both lower variance and negligible bias. While Ridge–REG attains a similar RMSE, this occurs at the cost of a larger bias—often exceeding that of REG—because the regularization distorts finite-sample unbiasedness. Moreover, the confidence intervals derived from our estimators, using the Huber–White variance estimator discussed in Sections˜3.5 and 4.4, exhibit coverage rates remarkably close to the nominal level, whereas both REG and Ridge–REG tend to undercover substantially. We note that the bias of LOORA–DM in these experiments is somewhat larger than that of LOORA–HT, because the treatment assignments are generated under simple random assignment rather than complete random assignment. These findings confirm that leave-one-out regularization provides a principled way to stabilize regression adjustment without sacrificing unbiasedness or inferential validity.

### v2

@@ 5.2 Comparison with Gram–Schmidt Random Walk Design
Harshaw et al. (2024) proposes an ingenious design-based approach to variance reduction. Rather than modifying the estimator, they optimize the randomization mechanism itself through a Gram-Schmidt Walk (GSW) design that ensures covariate balance while controlling the marginal assignment probabilities. One potential limitation of GSW is that it requires covariate information on all sample units prior to treatment assignment. In contrast, simple or complete random assignment can be implemented online (sequentially) as additional experimental subjects are recruited for a study. Moreover, with simple or complete randomization, a researcher can freely add or remove covariates, analyze different subgroups, or draw additional samples even after the experiment has been conducted. This flexibility, together with the transparency and ease of implementation of standard randomization, makes simple and complete randomization attractive in many applied settings. Motivated by these considerations, LOORA-HT and LOORA-DM are explicitly designed for simple and complete randomization, respectively.
Harshaw et al. (2024) shows that the GSW design achieves the following variance bound for any choice of $\phi\in(0,1)$:
$\frac{1}{n^{2}}\min_{\boldsymbol{\beta}\in\mathbb{R}^{k}}\Big[\frac{1}{\phi}\left\|{\boldsymbol{\mu}}-{\bm{X}}\boldsymbol{\beta}\right\|_{2}^{2}+\frac{\zeta_{{\bm{X}}}^{2}}{1-\phi}\left\|\boldsymbol{\beta}\right\|_{2}^{2}\Big].$ (13)
We now compare this bound with the LOORA-HT variance. The second term in (13) and the corresponding regularization term in (5) are not directly comparable in general, although both scale on the order of $k/n^{2}$. For the first term, with a constant probability of assignment,
$\min_{\boldsymbol{\beta}\in\mathbb{R}^{k}}\Big[\frac{1}{\phi}\left\|{\boldsymbol{\mu}}-{\bm{X}}\boldsymbol{\beta}\right\|_{2}^{2}+\frac{\zeta_{{\bm{X}}}^{2}}{1-\phi}\left\|\boldsymbol{\beta}\right\|_{2}^{2}\Big]=\min_{\boldsymbol{\beta}\in\mathbb{R}^{k}}\Big[\frac{1}{\phi}\left\|{\boldsymbol{\mu}}-\widetilde{{\bm{X}}}\boldsymbol{\beta}\right\|_{2}^{2}+\frac{\zeta_{\widetilde{{\bm{X}}}}^{2}}{1-\phi}\left\|\boldsymbol{\beta}\right\|_{2}^{2}\Big].$
Therefore, setting $\lambda=\zeta_{\widetilde{{\bm{X}}}}^{2}/(1-\phi)$, Lemma˜1 implies that
$\widetilde{h}_{\lambda ii}\;\leq\;\frac{1-\phi}{2-\phi}.$
@@
$\frac{1}{(1-\widetilde{h}_{\lambda ii})^{2}}\;\leq\;(2-\phi)^{2}.$
For $\phi\in[0,(3-\sqrt{5})/2)$, we have $(2-\phi)^{2}\leq 1/\phi$. Hence, the scaling factor that appears in our variance bound is uniformly smaller than that of the GSW design. Moreover, the first term in (13) diverges as $\phi\to 0$, whereas the corresponding constant in the variance of LOORA-HT remains bounded.
Finally, when the leverage scores of $\widetilde{{\bm{X}}}$ are small, LOORA can set $\lambda=0$. In contrast, under the GSW design the smallest feasible regularization parameter (i.e., coefficient of $\left\|\boldsymbol{\beta}\right\|_{2}^{2}$ in (13)) is $\lambda=\zeta_{{\bm{X}}}^{2}$.
5.3 Leave-One-Out and Leave-Two-Out Estimators
Aronow and Middleton (2013) studies a general class of unbiased Horvitz-Thompson estimators. These estimators adjust unit $i$’s outcome by a function $f({\bm{x}}_{i},\boldsymbol{\theta}_{i})$, where $\boldsymbol{\theta}_{i}$ is a vector of parameters. The estimator is unbiased when $d_{i}$ is uncorrelated with $f({\bm{x}}_{i},\boldsymbol{\theta}_{i})$. A useful case arises when $\boldsymbol{\theta}_{i}$ depends only on outcomes $y_{j}$ whose treatment indicators $d_{j}$ are independent of $d_{i}$. Under independent treatment assignment, this condition suggests leave-one-out constructions. LOORA-HT follows this logic by setting $\boldsymbol{\theta}_{i}=\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$, as defined in Algorithm˜1.
Wu and Gagnon-Bartsch (2018) studies a special case of Aronow and Middleton (2013) in which the adjustment term for observation $i$ is constructed as a linear combination of two counterfactual outcomes for unit $i$ (obtained by any method of our choosing, such as linear regression or random forests)—a predicted outcome for unit $i$ under treatment and a predicted outcome under control—via a leave-one-out fit, which satisfies the sufficient condition for $f$ stated in Aronow and Middleton (2013).
Building on these contributions, Spiess (2025) provides a general characterization of unbiased regression-adjusted estimators for average treatment effects under arbitrary randomization schemes when the treatment-assignment probability is the same for all units. He shows that for Horvitz–Thompson and difference-in-means type estimators, unbiasedness imposes specific structural constraints on how each unit’s outcome can depend on treatment assignments of other units. In particular, under simple random assignment with equal treatment assignment probability for all units, the Horwitz–Thompson estimator must have a leave-one-out structure—each adjusted outcome for unit $i$ may depend on the treatment assignments of all other units except ii—whereas under complete random assignment, the estimator must satisfy a stronger leave-two-out condition, meaning that each pairwise contribution between units $i$ and $j$ can depend only on the treatment assignments of units other than $i$ and $j$.
We next show that LOORA-DM satisfies Spiess’s leave-two-out requirement and therefore meets the necessary conditions for unbiasedness in Spiess (2025). Specifically, we show that the LOORA–DM estimator admits the representation
@@ 5.3 Leave-One-Out and Leave-Two-Out Estimators
where $\phi_{ij}$ is an adjustment function and ${\bm{z}}_{-ij}$ collects the treatment assignments of all units other than $i$ and $j$. Note that
@@
We define
$\widetilde{y}_{\ell}=\begin{cases}\displaystyle\frac{n_{C}(n-1)}{(n_{T}-1)n}y^{(1)}_{\ell}&\mbox{if }d_{\ell}=1,\\[2.58334pt] \displaystyle\frac{n_{T}(n-1)}{(n_{C}-1)n}y^{(0)}_{\ell}&\mbox{if }d_{\ell}=0,\end{cases}$
@@
Some algebra shows that these expressions are exactly the adjustment values in Algorithm˜2 for both treated and control units. Thus, $\widehat{\tau}_{\textup{LDM}}$ admits a leave-two-out representation. Section˜B.3 gives a detailed proof.
6 Evidence from Within-Subject Experiments
This section presents empirical evidence on the finite-sample performance of LOORA estimators. The analysis uses experimental data from two within-subject studies. In a within-subject experiment, each unit experiences both the treatment and control conditions, so both potential outcomes are observed for every unit. We use within-subject experiment data to simulate randomized evaluations with known ground truth and with a realistic joint distribution of potential outcomes.
The datasets are drawn from Allcott and Taubinsky (2015) and McDonald and Hanmer (2025). A brief description of each is provided below.
@@ 6 Evidence from Within-Subject Experiments
    Statehood dataset. McDonald and Hanmer (2025) studies the persuasiveness of policy arguments using a within-subject design. Respondents first record their opinions on a policy issue, then read one or more arguments either supporting or opposing the policy, and finally re-state their opinions after exposure to each argument. We use the portion of their data concerning opinions about granting statehood to the District of Columbia (DC). The outcome under control is the respondent’s opinion before reading any arguments; the outcome under treatment is the opinion after reading (i) an argument against DC statehood emphasizing political corruption, and (ii) an argument in favor highlighting taxation without representation. We restrict attention to male respondents aged 40-49. Covariates include indicators for the different values of six categorical variables: party affiliation, race, voter registration, political attentiveness, education, and ideology. The resulting dataset contains $36$ units and $32$ covariates.
@@
    Lightbulb dataset. Allcott and Taubinsky (2015) conducts an experiment on consumer preferences between two types of lightbulbs offered at varying relative prices. Respondents first choose between the two options, after which they receive information about the energy costs of each type and make the choice again. The initial choice forms the control outcome; the post-information choice forms the treatment outcome. Covariates are binary indicators of categorical features from the original data, including employment status, renter status, U.S. region, living in a metropolitan area, marital status, income, housing type, gender, and ethnicity. The final dataset contains $123$ units and $53$ covariates.
Table 2: Simulation results for Statehood dataset (covariate-dependent assignment, $\alpha=0.05$).
ATE Estimators
OLS LOORA-HT
Linear Interacted $\lambda{=}\zeta^{2}$ $\lambda{=}2\zeta^{2}$ $\lambda{=}3\zeta^{2}$
Bias 0.045 0.047 0.000 0.000 0.000
SD 0.519 0.639 0.314 0.296 0.290
RMSE 0.521 0.640 0.314 0.296 0.290
CI Coverage (HC0) 0.727 0.167 0.948 0.950 0.951
CI Coverage (HC2) 0.929 0.723 0.948 0.950 0.951
Table 3: Simulation results for Lightbulb dataset (covariate-dependent assignment, $\alpha=0.05$).
ATE Estimators
OLS LOORA-HT
Linear Interacted $\lambda{=}\zeta^{2}$ $\lambda{=}2\zeta^{2}$ $\lambda{=}3\zeta^{2}$
Bias 0.002 0.001 0.000 0.000 0.000
SD 0.106 0.152 0.089 0.086 0.084
RMSE 0.106 0.152 0.089 0.086 0.084
CI Coverage (HC0) 0.911 0.715 0.956 0.958 0.958
CI Coverage (HC2) 0.960 0.942 0.955 0.957 0.958
We simulate randomized assignments by masking one outcome per unit according to a prescribed randomization rule. This setup lets us assess each estimator’s finite-sample performance, bias, standard deviation, RMSE, and the average coverage of confidence intervals, under realistic distributions for potential outcomes and covariates. We consider three treatment-assignment mechanisms. In the first two designs, treatment assignments are independent across units. In the first design, treatment assignment probabilities depend on covariate values. Specifically, we draw a random Gaussian vector in ${\mathbb{R}}^{k}$, compute for each unit $i$ the cosine similarity $c_{i}$ between this vector and its covariate vector, and set
@@
The second design is simple random assignment, with all treatment assignment probabilities equal to $0.5$. The third design is complete random assignment, with half the units assigned to treatment.
Simulation results are based on 100,000100{,}000 repetitions. Tables˜2 and 3 report results for covariate-dependent assignment; Tables˜4 and 5 for simple random assignment with equal probabilities; and Tables˜6 and 7 for complete random assignment. For simple random assignment, we compare our estimator, LOORA–HT, with two benchmarks: (i) linear regression-adjustment, and (ii) the regression adjustment of Lin (2013) with additional interacted terms. For complete random assignment, in addition to these, we also compare our estimator, LOORA–DM, with the exact bias correction approach of Chang et al. (2024), which we refer to as “Debiased” in Tables˜6 and 7.
We report the coverage probabilities of confidence intervals in Tables˜2, 3, 4, 5, 6 and 7 using heteroskedasticity-consistent (HC) variance estimators. HC0 denotes White’s original HC estimator (White, 1980), while HC2 denotes the leverage-adjusted HC estimator of MacKinnon and White (1985). The variance estimator in Theorem˜4 corresponds to HC0, whereas the variance estimator in Theorem˜7 corresponds to HC2.
All simulations use recentered covariates and an intercept. That is, regression adjustment uses the matrix of regressors
$\boldsymbol{M}=\begin{bmatrix}{\bm{X}}-\overline{{\bm{X}}}&~~\boldsymbol{1}\end{bmatrix}$.
The value of $\zeta$ in the tables is equal to $\left\|\boldsymbol{M}\right\|_{2,\infty}$.
Throughout Tables 27, LOORA estimators consistently achieve lower bias, higher precision, and lower RMSE than competing methods. Moreover, the confidence intervals for LOORA estimators based on heteroskedasticity-robust standard errors exhibit coverage rates remarkably close to the nominal level, whereas the alternative methods tend to undercover substantially. These results demonstrate the potential role of LOORA and ridge regularization to stabilize regression adjustment without sacrificing unbiasedness or inferential validity.
Table 4: Simulation results on Statehood dataset (simple randomization, $\alpha=0.05$).
ATE Estimators
OLS LOORA-HT
Linear Interacted $\lambda{=}\zeta^{2}$ $\lambda{=}2\zeta^{2}$ $\lambda{=}3\zeta^{2}$
Bias 0.047 0.048 0.000 0.000 0.000
SD 0.512 0.628 0.299 0.285 0.279
RMSE 0.514 0.630 0.299 0.285 0.279
CI Coverage (HC0) 0.724 0.149 0.952 0.954 0.955
CI Coverage (HC2) 0.927 0.720 0.953 0.953 0.955
Table 5: Simulation results on Lightbulb dataset (simple randomization, $\alpha=0.05$).
ATE Estimators
OLS LOORA-HT
Linear Interacted $\lambda{=}\zeta^{2}$ $\lambda{=}2\zeta^{2}$ $\lambda{=}3\zeta^{2}$
Bias 0.002 0.001 0.000 0.000 0.001
SD 0.105 0.134 0.084 0.082 0.081
RMSE 0.105 0.134 0.084 0.082 0.081
CI Coverage (HC0) 0.909 0.733 0.961 0.962 0.963
CI Coverage (HC2) 0.961 0.947 0.961 0.961 0.962
Table 6: Simulation results on Statehood dataset (complete randomization, $\alpha=0.05$).
ATE Estimators
OLS Debiased LOORA-DM
Linear Interacted Linear Interacted $\lambda{=}\zeta^{2}$ $\lambda{=}2\zeta^{2}$ $\lambda{=}3\zeta^{2}$
Bias 0.046 0.043 0.001 0.039 0.001 0.001 0.001
SD 0.501 0.615 0.311 0.410 0.282 0.274 0.272
RMSE 0.503 0.616 0.311 0.412 0.282 0.274 0.272
CI Coverage (HC0) 0.726 0.104 0.901 0.147 0.955 0.958 0.959
CI Coverage (HC2) 0.928 0.701 0.985 0.723 0.961 0.963 0.964
Table 7: Simulation results on Lightbulb dataset (complete randomization, $\alpha=0.05$).
ATE Estimators
OLS Debiased LOORA-DM
Linear Interacted Linear Interacted $\lambda{=}\zeta^{2}$ $\lambda{=}2\zeta^{2}$ $\lambda{=}3\zeta^{2}$
Bias 0.002 0.003 0.000 0.001 0.000 0.000 0.000
SD 0.105 0.128 0.085 0.089 0.082 0.080 0.080
RMSE 0.105 0.128 0.085 0.089 0.082 0.080 0.080
CI Coverage (HC0) 0.910 0.748 0.964 0.886 0.963 0.962 0.963
CI Coverage (HC2) 0.962 0.951 0.989 0.985 0.964 0.965 0.965